OpenStudy (anonymous):

Medal !! FAN !! plz which line is a graph of the equation ? -2x+5y=10? A.line a B.line b C.line c D.line d

OpenStudy (godlovesme):

graphs?

OpenStudy (anonymous):

first put the equation in a form more familiar such as slope - intercept form => y = mx+b so solve the equation for 'y'

OpenStudy (anonymous):

yes @Godlovesme

OpenStudy (anonymous):

y = (2/5)x +2

OpenStudy (anonymous):

if the value of m is positive what direction does the line go?

OpenStudy (godlovesme):

nmv @manumben is helping :)

OpenStudy (anonymous):

right @manumben ?

OpenStudy (anonymous):

as you move to the right does it go up or down?

OpenStudy (anonymous):



OpenStudy (anonymous):

up @manumben

OpenStudy (texaschic101):

-2x + 5y = 10 to find the y intercept, sub in 0 for x -2(0) + 5y = 10 5y = 10 y = 2 ....so the line crosses the y axis at (0,2) to find the x axis, sub in 0 for y -2x + 5(0) = 10 -2x = 10 x = -5 ... so the line crosses the x axis at (-5,0) so look at your points (0,2) and (-5,0) and see which line goes across these points

OpenStudy (anonymous):

yes and when the equation of a line is in 'slope intercept' form it comes from the fact the number next to 'x' is the slope and the number with out is where on the y-axis it will cross when 'x' = 0